Question #19 please. Topic is Exponential and Logarithmic Functions/Exponential Growth and Decay. The answers supposed to be 760 millimeters is the atmospheric pressure at sea level; 19i. 740.086 mm; 19ii. 535.753 mm; 19iii. 469.682 mm. Please show steps for these answers. I would like to see and understand it thoroughly. Thank you so much! Prelalu times per year: Use the compound interest formula to determine ind interest formula to determine A, the amount you St
Answer #1

Q.19. P = 760e-0.11445h where h is in km height above sea level

at sea level , h = 0

thus P(h = 0) = 760e-0.11445*0 = 760 mm of mercury

(i) at Diamond head crater i.e. h = 0.232 km , P(h = 0.232) = 760e-0.11445*0.232 = 740.086 mm of mercury

(ii) at Haleakala crater i.e. h = 3.055 km , P(h = 3.055) = 760e-0.11445*3.055 = 535.753 mm of mercury

(iii) at Maunakea obsevatory i.e. h = 4.205 km , P(h = 4.205) = 760e-0.11445*4.205 = 469.682 mm of mercury